# Understanding the Challenges and Promises of Developing Generative AI Apps: An Empirical Study

### v3

@@ 5.1. Experiment Setup
In RQ1, we experiment with varying levels of data cleaning and numbers of few-shot examples. We find that filtering out non-informative reviews using an LLM and providing five-shot examples improves topic extraction and assignment accuracy. In RQ2, we apply the optimal topic extraction prompt identified in RQ1 to the remaining app categories and assign topics to all reviews in Stage 3. In this section, we provide a detailed description of the implementation of RQ2.
Topic extraction and assignment. The results of RQ1 indicate that the optimal dataset for analysis is Stage 3 data, which underwent four stages of cleaning: basic cleaning, temporal filtering, short review exclusion, and LLM-assisted informative review filtration. Our dataset spans 11 app categories, totaling 676,066 user reviews as shown in Table 2. As in RQ1, we sample a statistically significant set of informative reviews (large sample) per app category, using a 95% confidence level and 2% margin of error. The resulting sample sizes range from 761 to 2,368 reviews, with an average of 2,031 per category. For each app category $y$, we extract the top 10 topics from the large sample using the validated 5-shot topic extraction prompt shown in Figure 5. Next, we use the topic assignment prompt shown in Figure 6 to assign one of the extracted topics to each review in the dataset. For each app category $y$, the corresponding top 10 topics are used as input to label reviews with a single topic.
Classification and categorization. As our primary interest lies in understanding user feedback related to Gen-AI functionality, we identify and analyze reviews that discuss such features. To do so, we manually examine all extracted topics and label them as either Gen-AI topics, which reference Gen-AI capabilities, or non-Gen-AI topics, which pertain to general app features outside the scope of Gen-AI. All reviews originate from Gen-AI apps; the distinction here is between review types, not between Gen-AI and non-Gen-AI apps. We further group semantically related topics into high-level topic categories based on shared themes or functionality, facilitating higher-level interpretation of user concerns. In line with prior studies (Maram_LLM_Cure; Pagano_Maalej_2013; Chen_Lin_Hoi_Xiao_Zhang_2014), we also classify reviews with 1 or 2-star ratings as negative, 3-star rating as neutral, and 4 or 5-star rating as positive.
Analysis. We perform large-scale quantitative analysis on all automatically labeled reviews across both app categories and topic categories, followed by qualitative analysis on a representative sample. Specifically, for the quantitative component, we compute the following metrics:
@@
    Average Gen-AI Rating (AGR): the average star rating of reviews labeled with Gen-AI topics.
@@
    Average Non-Gen-AI Rating (ANR): the average rating of reviews labeled with non-Gen-AI topics.
@@
We conduct the Mann–Whitney U test to examine whether reviews labeled as Gen-AI-related receive significantly higher ratings than those that are not. For developer-response analysis, we compute reply ratios, response delays, and response-type distributions for both Gen-AI and non–Gen-AI reviews, and we compare per-app reply behavior using a Wilcoxon signed-rank test (wilcoxon1992).
We complement our quantitative analyses with qualitative examination by manually coding statistically significant samples (95% confidence level, 10% margin of error) of both user reviews and developer responses. For user reviews we draw samples from every Gen-AI topic category resulting in a total of 901 reviews. We manually code and analyze all sampled reviews to provide richer, contextualized insights that may surface actionable implications for both researchers and developers.
For developer responses, we perform two qualitative analyses. First, we code a representative sample of 96 replies drawn across all app categories to characterize the overall distribution of developer responses. Second, we manually code an additional 281 replies from three focal categories (Games, Education, and Entertainment) to capture deeper, category-specific response practices. Together, these qualitative analyses provide contextual grounding for the quantitative findings and offer a more complete understanding of both user concerns and developer engagement within Gen-AI apps.
@@ 5.2. RQ2 Findings
Observation 2.1: Gen-AI reviews consistently receive higher user ratings (AGR) than non-Gen-AI reviews (ANR) across all app categories of Gen-AI apps. As shown in Table 5, the average rating of reviews discussing Gen-AI functionality (i.e., AI Performance and Utility & Use Cases) exceeds the average rating of non-Gen-AI reviews (i.e., Technical Difficulties and Monetization Methods & Structure) in every app category of Gen-AI apps. This indicates that, when users explicitly comment on Gen-AI features, they tend to rate their experiences more positively than when they comment on other aspects of the same apps.
To assess the statistical significance of this difference, we apply the Mann–Whitney U test. The results reveal a significant difference in the rating distributions between the two groups (U=65,113,513,850U=65{,}113{,}513{,}850, $p<0.0001$). Gen-AI reviews tend to receive higher ratings, with a median rating of 5.0 and a mean of 4.10, compared to a median of 4.0 and a mean of 3.32 for non-Gen-AI reviews. These findings suggest that Gen-AI functionality is a major contributor to positive user experiences across a wide range of app categories.
This pattern is especially pronounced in categories such as Education and Books & Reference, which show some of the largest gaps between AGR and ANR. In Books & Reference, Gen-AI topics such as AI Performance and Human AI Interaction & Emotional Connection receive high ratings, while non-Gen-AI topics like Technical Difficulties and Customer Support are rated much lower due to crashes, generation failures, and unresolved support issues. Similarly, in Education, users highly rate Gen-AI topics such as Content Quality and Utility & Use Cases, but express dissatisfaction with Monetization Methods & Structure and Technical Difficulties. Together, these results indicate that strong Gen-AI features can substantially boost user satisfaction, but poor reliability, monetization, or support can still pull overall ratings down. Developers should therefore treat Gen-AI capabilities as amplifiers of perceived value, not as substitutes for robust app quality, stability, and fair pricing.
@@
Productivity 4.2 3.5 84,122 50% 6
Entertainment 4.0 3.2 56,672 48% 6
Education 4.6 3.3 31,465 75% 6
Art & Design 3.4 3.2 26,585 21% 3
Photography 3.9 3.3 15,567 12% 2
Tools 4.0 3.4 7,782 22% 3
Health & Fitness 4.8 4.0 6,896 88% 7
Video Players & Editors 4.6 3.9 5,179 14% 2
Games 3.6 2.7 1,788 35% 5
Music & Audio 3.0 2.6 1,679 18% 4
Books & Reference 4.1 2.3 428 39% 4
Average 4.0 3.2 21,651 38% 4
Observation 2.2: Gen-AI functionality is a prominent focus in user feedback, comprising an average of 38% of reviews across app categories and reaching as high as 88% for the Health & Fitness category. As shown in Table 5, the prevalence of Gen-AI–focused reviews varies by app category, reflecting the differing roles Gen-AI plays across app domains. Categories such as Health & Fitness, Education, and Productivity contain the highest proportions of Gen-AI reviews, with 88%, 75%, and 50%, respectively, indicating that users in these domains directly engage with AI features to achieve personal or task-oriented goals such as self-improvement, learning, or workflow assistance.
@@
In contrast, categories such as Photography and Video Players & Editors exhibit lower percentages of Gen-AI reviews, where 58% of feedback still focuses on non-Gen-AI topics such as usability, technical issues, and subscription or pricing models. This suggests that in creative or media-focused apps, users remain more concerned with traditional performance and value-for-money factors than with AI-driven functionality.
Figure 7 illustrates the distribution of Gen-AI review percentages across individual apps within various categories, highlighting how consistently users discuss Gen-AI features within each domain. Categories such as Entertainment and Health & Fitness exhibit the widest spreads, suggesting substantial variation among apps; some are highly AI-centric, while others receive minimal user attention to Gen-AI features. In contrast, categories like Video Players & Editors and Photography show narrow spreads concentrated at low percentages, indicating that across these domains, Gen-AI features consistently play an assistive or complementary role rather than being the central focus of user experience. Together, these patterns suggest that while certain app categories demonstrate uneven user emphasis on Gen-AI, others achieve relatively uniform user engagement of Gen-AI features.
Observation 2.3: Users discuss a wide range of Gen-AI topics, revealing where developers excel and where improvement is needed. Table 6 and Table 7 present the top Gen-AI and non-Gen-AI topics, respectively. Developers should pay particular attention to topics that appear most frequently across apps because they represent the aspects users discuss and evaluate most. Excelling in these areas is critical for maintaining competitiveness and user satisfaction. The Gen-AI topics with the highest numbers of reviews are AI Performance, Utility & Use Cases, and Content Quality. The topics most commonly mentioned across different apps are AI Performance, Content Quality, and Content Policy & Censorship. Similarly, topics with high average ratings (e.g., Human AI Interaction & Emotional Connection and Utility & use cases) serve as benchmarks for best user experiences that other apps can emulate. Conversely, lower-rated or frequently criticized topics (e.g., Content Policy & Censorship and Content Quality) highlight recurring weaknesses across Gen-AI apps. Developers should treat these as priority areas for refinement to stay ahead of the competition and sustain user trust. We further examine each of the Gen-AI categories in the subsections below.
Gen-AI Topic Category GRC AGR # Apps # App Categories
AI Performance 61,071 4.0 127 10
Utility & Use Cases 56,505 4.8 45 3
Content Quality 44,594 3.6 131 8
Content Policy & Censorship 27,521 2.9 76 5
Human AI Interaction & Emotional Connection 25,142 4.8 13 4
Features & Functionality 16,275 4.0 58 4
Customer Support & Community 3,361 4.7 12 3
Comparison to Other Apps 2,315 3.9 13 3
Personalization & Customization 1,283 4.0 30 3
Accessibility & Inclusivity 96 4.1 4 1
Non-Gen-AI Topic Category NRC ANR # Apps # App Categories
Technical Difficulties 99,621 2.2 157 10
Monetization Methods & Structure 85,756 2.3 159 11
User Interface & Experience 80,827 4.5 156 11
Utility & Use Cases 21,464 4.8 34 1
Updates & Evolution 21,100 2.8 92 7
Personalization & Customization 6,860 4.8 18 1
Customer Support & Community 6,848 1.8 85 6
Limitations of Usage 5,971 2.5 45 4
Features & Functionality 1,810 3.0 11 1
Comparison to Other Apps 1,493 3.4 17 1
Accessibility & Inclusivity 401 2.6 5 2
Other (unassigned) 105,752 4.2 165 11
1) AI Performance. AI Performance refers to how well the AI model performs in terms of 1) accuracy (27%), 2) speed (22%), 3) understanding (15%), 4) memory (5%), 5) reliability (5%), and 6) creativity & randomness (3%). AI accuracy describes the precision of the AI response to the user’s prompt. AI speed represents how quickly the AI responds to a user’s prompt. AI understanding indicates the AI’s ability to comprehend the meaning, intent, and context behind the user’s prompt. AI memory signifies how well the AI remembers prior conversations and links them to the current one. AI can have both short-term memory, which spans a single conversation, and long-term memory, which spans multiple conversations. Reliability refers to the consistency and dependability of the AI system in performing its intended tasks without crashing, freezing, timing out, or failing to generate output. Creativity captures the degree of originality, uniqueness, or unpredictability in the AI’s responses.
Figure 8. Examples of reviews discussing AI Performance
Insights. We find that 79% of the users who discussed AI Performance provided a positive review with a rating of 4 or 5. Reviews 1 and 2 in Figure 8 show examples of positive reviews. Despite this overall satisfaction, 16% of reviews mentioned limitations in AI Performance, with the most frequently cited issue being AI understanding (5%), followed by inconsistency in accuracy (4%), reliability (3%), speed (2%), and AI memory (1%).
A recurring theme in the less favourable reviews is poor AI understanding, which leads to irrelevant or partially correct responses. Some users show awareness of prompt engineering and attempt to guide the AI with examples, yet still encounter issues. For instance, review 3 in Figure 8 illustrates a user struggling to communicate effectively with the AI, despite providing clear examples. Similarly, review 4 in Figure 8 highlights inconsistent performance for complex requests. While the AI excels at handling simple tasks, it often struggles with more complex ones.
Although memory-related concerns (e.g., forgetting prior context or repeating content) are occasionally mentioned as shown in review 5 in Figure 8, they tend to be rated more leniently, indicating that users may be more forgiving of these limitations.
@@
2) Content Quality. Content Quality refers to users’ assessment of AI-generated content in terms of usefulness, coherence, creativity, engagement, structure and presentation.
Figure 9. Examples of reviews discussing Content Quality
Insights. In our sample, 67% of reviews express satisfaction, praising the creative and polished outputs such as producing “radio-ready” songs and appending useful resources to textual answers as illustrated in reviews 1 and 2 in Figure 9 respectively.
However, 33% of reviews reveal concerns in three domains. First, users report that image-generation features frequently fail to produce accurate or faithful visual outputs, particularly for complex or creative prompts (e.g., failed attempts to depict two people and an eagle wearing denim shorts in reviews 3 and 4 in Figure 9, respectively.) This finding aligns with broader evidence that AI image generators often collapse under high-complexity tasks (Petsiuk_Siemenn_Surbehera_Chin_Tyser_Hunter_Raghavan_Hicke_Plummer_Kerret_et_al_2022). Second, 8% of reviews highlight concerns regarding insufficient diversity and inclusion. For instance, review 5 in Figure 9 highlights a disparity in image quality between indigenous women and people of colour. Third, a minority of users (3% of reviews) highlight formatting and user interface issues in text outputs, such as grammatical errors, excessive length, inconsistent structure, indicating a need for improved polish as illustrated in review 6 in Figure 9.
$\Rightarrow$ Implication 1.
Gen-AI reviews consistently yield higher user ratings across app categories. Developers must recognize Gen-AI as a crucial driver of overall user satisfaction and strategically incorporate these features to remain competitive. However, they should treat Gen-AI capabilities as amplifiers of perceived value, not as substitutes for robust app quality, stability, and fair pricing. Even the most advanced AI functionality cannot compensate for poor technical reliability, weak customer support, or frustrating monetization structures. While users are generally satisfied with the overall performance of Gen-AI apps, a portion of their concerns relate to limitations in human-AI communication, particularly when the AI struggles to interpret nuanced or complex prompts. To better meet user expectations and improve usability, developers should prioritize enhancing the Gen-AI’s ability to understand and respond appropriately to more complex interactions. Although users generally regard Gen-AI outputs as high quality, enhancing image accuracy, diversity, inclusion, formatting, and user interface presentation remains essential for delivering truly reliable and user-centered content. Developers should prioritize these areas to refine trust, inclusivity, and satisfaction across varied use cases.
@@
3) Content Policy & Censorship. Content policy refers to rules and guidelines that define what type of AI-generated content is permissible. Censorship refers to automated or manual mechanisms used to block, modify, or suppress inappropriate, offensive, harmful, or illegal content based on those rules.
Figure 10. Examples of reviews discussing Content Policy & Censorship
Insights. In our review sample, 73% of users comment on censorship experience. Of these 23% report frustration with the filtering that often block benign content, adversely affecting usability. For example, the author of review 1 in Figure 10 reports getting censored for saying “hi.” On the other hand, 10% of the raised concerns relate to lack of effective filtering such as unethical imagery involving children bypassing filters as expressed in review 2 in Figure 10. Regarding adult content, 67% of concerned users express dissatisfaction with the excessive restriction of access. While 19% explicitly request an optional censorship toggle for adult users as illustrated in review 3 in Figure 10. Additionally, users express concern over unclear content ownership. Some users report that their creations were reused or uploaded by the platform without consent, as shown in review 4 in Figure 10.
$\Rightarrow$ Implication 2.
Developers need to strike the right balance in filtering content to avoid blocking genuinely harmful material while preventing harmless content. Ideally, developers could offer optional and adjustable censorship settings for mature or context-specific use. At the same time, platforms must clearly state who owns AI-generated outputs and how they can be reused, to maintain transparency and protect user trust.
4) Features & Functionality. Features & Functionality refer to the Gen-AI-driven capabilities offered by the app, such as voice interaction and prompt autocompletion, which shape the user experience and determine which features users find valuable or lacking.
Insights. In our sample, 75% of user reviews praise specific Gen-AI features, while 25% criticize them. For example, some users react negatively when a “stop-generation” button is removed and replaced with a voice mode feature, as illustrated in review 1 in Figure 11. While 11% of users enjoy the voice features, some describe them as useless. This indicates that even well‑received Gen-AI additions (such as voice interaction) can alienate users if they disrupt existing valued controls.
Figure 11. Examples of reviews discussing Features & Functionality
5) Utility & Use Cases. Utility & Use Cases refer to the practical roles Gen-AI plays in users’ daily lives, including enhancing accessibility, supporting professional tasks, and enabling creative expression across diverse contexts.
Figure 12. Examples of reviews discussing Utility & Use Cases
Insights. In our sample, 64% of reviews emphasize AI’s educational utility, such as helping with homework, essay writing, tutoring, and breaking down complex topics as shown in review 1 in Figure 12. Additionally, some users specifically praise how Gen-AI can act as an accessible on-demand learning assistant for dyslexic users, as commented in review 2 in Figure 12. Beyond education, 11% of reviews describe how the Gen-AI tools facilitate routine work tasks, such as inspiring new ideas, structuring project workflows, and even acting as fitness or guitar instructors, as illustrated in review 3 in Figure 12.
6) Personalization & Customization. Personalization & Customization describe how Gen-AI systems adapt to user preferences and allow users to tailor outputs. This includes customizing output formats, editing generated content, and adopting preferred AI personas, e.g., voices or AI character roles.
Figure 13. Examples of reviews discussing Personalization & Customization
Insights. In our sample, 38% of reviews emphasize the AI’s inability to create a personalized experience. For example, review 1 in Figure 13, discusses the app’s inability to maintain shopping and to-do lists. Additionally, 10% of users request inline editing capabilities to support customization by refining generated content directly, rather than having to regenerate entirely. For instance, review 2 in Figure 13 suggests the option to edit lyrics in an already generated song rather than regenerate a new one entirely. Meanwhile, 9% of reviews request more diversity in AI characters and custom voices. This reflects a growing interest in personalizing AI identities to match preferences such as female voice options. For example, review 3 in Figure 13 describes the absence of customized voice features as a “deal breaker.”
$\Rightarrow$ Implication 3.
Developers should design Gen-AI features, such as voice mode, in a way that complements existing controls rather than replacing them. For instance, it is crucial to ensure new AI-driven tools coexist with familiar functionalities, so that developers can both innovate and maintain intuitive user experiences. Developers should implement customizable output options, enabling users to format, export, and save AI-generated content (e.g., lists, summaries) in a manner that suits their workflows. They should also add inline editing tools, so users can refine content directly without needing to regenerate responses. Additionally, offering persona customization, such as selectable character roles or voice styles, will enhance user engagement and make interactions feel more personalized. Despite the launch of tools like ChatGPT Edu (openaiChatGPTedu), which is tailored for academic contexts, Gen-AI platforms must go further by prioritizing accessibility-first design, for example, enabling assistive note-taking, voice transcription, and dyslexia support features. Developers should also integrate customizable, role-based modes or templates, such as “legal advisor,” “fitness coach,” or “creative brainstorming partner,” to make these tools more adaptable and effective across diverse professional and personal applications.

### v4

@@ 5.1. Experiment Setup
We apply RQ1’s optimal configuration to identify and analyze topics from user reviews. This section describes the topic extraction process, the classification of extracted topics, and the quantitative and qualitative analyses used to interpret user feedback. We implement our framework to both Google Play Store and App Store reviews and discuss platform comparison in Section 7. In this section, our analysis is focused on the Google Play Store dataset because it is larger and more substantive.
Topic extraction and assignment. Following RQ1, we draw a statistically significant large sample of informative reviews from stage 3 (2,291 average reviews per app categories) to extract the top 10 topics from each app category. These topics are then used to assign a single topic label to each review in the full dataset.
Classification and categorization. Our primary interest is user feedback related to Gen-AI functionality. We therefore manually examine the extracted topics and label them as either Gen-AI topics, which reference Gen-AI capabilities, or non-Gen-AI topics, which refer to general app functionality. All reviews originate from Gen-AI apps; the distinction therefore concerns the type of feedback rather than the app type. To facilitate higher-level interpretation, we group semantically related topics into broader topic categories. In line with prior studies (Assi et al., 2025a; Pagano and Maalej, 2013; Chen et al., 2014), we also categorize reviews by star ratings: 1-2 stars as negative, 3 stars as neutral, and 4-5 stars as positive.
Analysis. We perform large-scale quantitative analysis on all labeled reviews across both app categories (e.g., Productivity and Photography) and topic categories (e.g., AI Performance and Creative Potential). Specifically, we compute the following metrics:
@@
    Average Gen-AI Rating (AGR): the average rating of reviews discussing Gen-AI topics.
@@
    Average Non-Gen-AI Rating (ANR): the average rating of reviews discussing non-Gen-AI topics.
@@
To examine whether reviews discussing Gen-AI functionality receive higher ratings than other reviews, we compute AGR and ANR for each app by aggregating review scores at the app level. We apply the Wilcoxon signed-rank test (Wilcoxon, 1992) to compare the two ratings. To complement the quantitative results, we conduct qualitative coding of statistically significant samples (95% confidence level, 10% margin of error). We sample reviews from each Gen-AI topic category, resulting in 762 reviews that are manually coded to identify recurring user concerns that derive opportunities, challenges, and actionable insights.
@@ 5.2. RQ2 Findings
Finding 2.1: Gen-AI reviews consistently receive higher ratings than non-Gen-AI reviews across app categories, suggesting that negative feedback often relates to app design issues rather than the Gen-AI functionality itself. Table 5 shows that reviews discussing Gen-AI functionality receive higher ratings than those discussing other app aspects, with an average of 4.2 compared to 2.8. A Wilcoxon signed-rank test confirms that AGR is significantly higher than ANR across apps ($p<0.001$), with significant differences observed in 7 of 9 categories. This pattern is particularly pronounced in categories such as Art & Design and Music & Audio, where Gen-AI topics related to Creative Potential receive very high ratings (4.8-4.9), while non-Gen-AI topics such as Monetization Methods & Structure receive substantially lower ratings (1.8).
@@
Productivity 4.3 3.0 257,434 64% 5
Photography 4.5 2.8 66,066 39% 3
Entertainment 3.6 2.8 61,803 41% 5
Art & Design 4.3 2.3 54,227 43% 3
Education 4.5 4.0 46,942 72% 6
Games 3.9 2.4 15,695 62% 6
Tools 4.3 3.0 14,923 41% 5
Video Players & Editors 4.5 3.2 12,741 32% 3
Music & Audio 4.1 2.1 6,992 39% 5
Average 4.2 2.8 59,647 48% 5
These results suggest that lower ratings are often driven by non-Gen-AI concerns, particularly monetization and technical issues, rather than dissatisfaction with Gen-AI functionality itself. For developers, these results suggest that strong Gen-AI capabilities alone are insufficient, as poor app design, reliability issues, and monetization practices can override their benefits.
Finding 2.2: The percentage of reviews discussing Gen-AI varies across app categories, reflecting whether Gen-AI is a core or supplementary feature. Table 5 shows variation in the proportion of Gen-AI-related reviews across categories. Categories such as Education (72%) and Productivity (64%) exhibit the highest percentages, indicating that users primarily evaluate these apps through their AI capabilities. In these domains, discussions frequently center on AI performance and practical use cases, such as studying, writing, and content generation, highlighting strong user interest in the applicability of Gen-AI features.
In contrast, categories such as Video Players & Editors, Music & Audio, and Photography show much lower proportions of Gen-AI reviews, where AI functionality typically serves as a supplementary feature within broader tools. As a result, user feedback is more focused on non-Gen-AI concerns such as technical issues, monetization, and usability.
Figure 4 further shows variation within categories, indicating that apps differ in how they position Gen-AI functionality. For developers, these results suggest that when Gen-AI is a core feature, users focus on its utility and performance, whereas when it is supplementary, overall app design and reliability play a larger role in shaping user perceptions.
@@
Finding 2.3: User discussions of Gen-AI apps reveal key opportunities, including AI for Accessibility and Wellbeing, AI as a Collaborative Creative Tool, and AI Versatility, as well as challenges related to Managing User Expectations and AI Limitations, Balancing Content Moderation and Creative Freedom, and Strategic Integration of Gen-AI Features. Table 6 lists the Gen-AI topics with their descriptions and examples. Descriptive statistics on Gen-AI and non-Gen-AI topic categories can be found in Tables 12 and  13 in Appendix C. Building on the topic landscape identified, user feedback reveals several recurring themes that highlight both the potential and the limitations of Gen-AI apps. We synthesize these themes into three key opportunities and three challenges for developers designing Gen-AI apps as follows.
Opportunity 1: Discussions of accessibility and wellbeing in Gen-AI apps highlight opportunities for inclusive and responsible design with appropriate safeguards for vulnerable users. User discussions reveal that Gen-AI apps support accessibility needs, learning assistance, and emotional wellbeing. These use cases are primarily reflected in the Utility & Use Cases and Emotional Connection topics, which receive some of the highest average ratings in our dataset (4.8).
Reviews in the Utility & Use Cases topic show that users rely on Gen-AI tools for educational and everyday assistance. In our sample, 59% of reviews discussing this topic describe educational use cases such as tutoring, explaining difficult concepts, or assisting with homework. Users also highlight accessibility benefits, describing how Gen-AI systems help individuals with learning disabilities such as dyslexia or Attention-deficit/hyperactivity disorder (ADHD) by providing personalized explanations and on-demand assistance.
Similarly, the Emotional Connection topic reveals that users frequently interact with Gen-AI systems as conversational companions. Reviews in this category are overwhelmingly positive (95%), with users describing their interaction with AI as a form of companionship, including having someone to talk to (21%), forming friendly relationships with the AI (6%), or using the system as a romantic companion (11%). Users report that interacting with the AI helps alleviate loneliness (14%), boredom (10%), or mental health concerns like anxiety and depression (5%). Although relatively rare, a small number of reviews (6%) raise concerns about potentially unhealthy interactions. For example, users describe the experience of speaking with AI characters as addictive or isolating, reporting that extended interactions reduced their engagement with real-world activities and relationships.
Together, these discussions illustrate how Gen-AI apps extend beyond productivity tools to support accessibility and wellbeing. For developers, these findings highlight the opportunity to design Gen-AI systems that intentionally support educational accessibility and responsible conversational companionship while ensuring appropriate safeguards for vulnerable users. For policymakers and regulators, our findings highlight the need to examine the formation of human-AI relationships and to establish guidelines that address risks of over-attachment and dependency, particularly among vulnerable populations such as individuals with disabilities or mental health concerns.
Opportunity 2: Discussions of Gen-AI as a collaborative creative tool highlight opportunities to design systems that support iterative co-creation and human-AI collaboration. User reviews describe Gen-AI systems as collaborative partners that assist users in ideation and creative exploration rather than fully automating tasks. This opportunity is primarily reflected in the Creative Potential and Utility & Use Cases topics.
The Creative Potential topic receives overwhelmingly positive sentiment, with 98% of reviews expressing positive ratings. Users describe employing Gen-AI tools for a wide range of creative activities including graphic design, storytelling, roleplaying, music creation, and social media content production. For example, users report generating short videos for platforms such as TikTok or Instagram, experimenting with character creation, and using AI-generated content as inspiration for creative projects.
Evidence of collaborative workflows also emerges in the Utility & Use Cases topic. Within this topic, reviews describe users working alongside AI to support tasks such as idea generation (4%), research assistance (4%), cooking guidance (1%), and programming support (1%). In these cases, the AI does not complete the entire task independently; instead, users interact iteratively with the system, refining prompts and incorporating AI-generated suggestions into their own work.
These discussions suggest that users increasingly perceive Gen-AI systems as collaborative partners that augment human creativity and problem solving rather than replacing human effort entirely. For developers, these findings highlight the importance of designing Gen-AI systems that support iterative interaction, ideation support, and flexible co-creation workflows where users and AI contribute jointly to task completion.
Gen-AI Topic Category Description (D)
Example (E)
AI Performance D: AI Performance captures user feedback on how effectively the AI generates responses, including aspects such as accuracy, speed, conversational quality, versatility across tasks in different domains, memory, and understanding.
E: “Nice art but not really responding to my text well. It shows basic pictures not the complex combination I asked for so not that worth it for some more crazy experimentation ok for basic generation I guess.”
Utility & Use Cases D: Utility & Use cases refers to reviews discussing the different ways they utilize Gen-AI in their lives.
E: “ChatGPT is my new personal trainer and guitar teacher.”
@@
Content Quality D: Content Quality captures users’ assessments of the usefulness, coherence, creativity, engagement, structure, and presentation quality of AI-generated outputs.
E: “Great pictures just not what I asked for. I have tried 7 times and not one was correct. Meaning I ask for two people with specific details to clothes etc. I get either one great drawing of one person or a distorted drawing of two people but only three legs Hilarious but I was not aiming for humor”
Creative Potential D: Creative Potential captures how users leverage Gen-AI tools for creative expression, including content creation, design, storytelling, and interactive conversations.
E: “Great for TikTok and Instagram short videos. ”
@@
Content Policy & Censorship D: Content policy refers to rules and guidelines governing the use of training and user data in Gen-AI systems, including how data is collected, used, and regulated. Censorship refers to automated or manual mechanisms used to block, modify, or suppress inappropriate, offensive, harmful, or illegal content based on those policies.
E: “I’ve been having this app for almost a year now and I was enjoying it a lot. But recently whenever I’ve been having the most normal conversations with the AI bots it always gives me the info that somehow the messages don’t meet the guideline or something like that. I don’t know why it does that but I’m pretty disappointed. Just to clarify it again my conversations are absolutely normal.”
Features & Functionality D: Features & Functionality captures user feedback on specific AI-enabled capabilities offered within the app, particularly voice interaction features and their usability.
E: “Can’t even set an alarm by voice.”
Emotional Connection D: Emotional Connection captures user experiences where Gen-AI systems function as conversational companions, sources of emotional support, or social interaction partners.
E: “This app shouldn’t exist at all. It ruins lifes harms the environment. the addiction is a real thing. At first it’s fun but the more you use it the more it pulls you in to the point where you can’t put it down and I wonder how many people struggle with this in silence. It’s disgusting that you profit off of people’s well being.”